OpenStudy (anonymous):

(1+2i)(2-3i)/(2-i)(3+2i) how do I solve this? :)

OpenStudy (amistre64):

what do you mean by 'solve'?

OpenStudy (amoodarya):

(1+2i)(2-3i)/(2-i)(3+2i) =(2-3i+4i-6i^2)/(6-3i+2i-2i^2) =(2+6+i)/(6+2-i) =(8+i)/(8-i) now multiply by (8+i)/(8+i) then simplify

OpenStudy (anonymous):

See the attachment.